# Demystifying Language Model Forgetting with Low-rank Example Associations

### v6

@@ Appendix F Effect of Targeted Replay on New-Task Learning and Downstream Performance
Appendix G Towards Interpreting Fine-Grained Associations
@@
Semantic meanings of $k$-th component in the low-rank approximation of the association matrix $Z$. We perform further analysis into the patterns captured by the $k$-th singular value and singular vectors by identifying the most relevant learned tasks and upstream example domain to the component. For each $k$ and its corresponding component $\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$, we extract top 3 rows with the highest mean (i.e., top 3 relevant learned tasks $T_{i}$). We also extract top 50 columns with highest mean (i.e. top 50 relevant upstream examples) and the domain where these upstream examples are drawn from. For OLMo models, the domains are one of C4, common-crawl, Gutenberg books, Reddit, Science, StackOverFlow, and Wikipedia. We compare the distribution of domains in the top 50 upstream examples, and perform a $z$-test to determine upstream example domain that is significantly more or less forgotten compared to a prior domain distribution of top 50 most forgotten upstream examples (columns with highest mean in $Z$). The results are summarized in Table 8.
@@ Appendix H Effect of Model Training Setups on Forgetting
Low-rank approximations hold at various learning rates, batch sizes, and LoRA fine-tuning. Table 9 summarizes $R^{2}$ of low-rank approximations. Reducing learning rate from the default 2e-6 to 1e-6 causes very minor change of $R^{2}$; Nevertheless, we notice increased $R^{2}$ under a larger learning rate like 5e-6, indicating even simpler associations between the learned tasks and forgotten upstream examples. Under this overly large learning rate, the model suffers substantially more forgetting on all examples. Besides, increasing the batch size (under the same number of parameter update steps) causes an increase in $R^{2}$, indicating simpler associations, possibly due to reduced variance of learned models under a larger batch size. LoRA fine-tuning Hu et al. (2022) results in simpler associations despite the model forgets less than full parameter fine-tuning.
@@
Replay can change the patterns of forgetting compared to no replay, but the change is mild. We compute sample Pearson correlation coefficient between upstream example forgetting collected under different replay strategies, and more specifically (1) no replay (2) random replay and (3) MF-Offline replay on OLMo-7B and 8 Dolly tasks. We evaluate forgetting of upstream examples held-out from replay. We include correlations between upstream example forgetting under two different training random seeds as the baseline.
We summarize the results in Table 10. The correlations between no replay and replay are around 0.8, lower than the 0.94 baseline, but are still strongly positive. Replay example selection strategies (random or MF-offline) do not have a strong impact on the correlation. The results imply replay changes patterns of forgetting to a mild degree. Future works can study the limit where the patterns of forgetting experience notable changes.
@@ Appendix I Combining Embedding and Matrix Completion based Prediction
We consider the embedding based approaches better capture fine-grained knowledge conflicts that leads to forgetting of one example by learning the other (e.g., two documents stating contradictory facts) that is missed by matrix completion approaches that rely on statistics of forgetting.
Appendix J Synthetic Dataset Experiments
@@
To summarize, our analysis with synthetic datasets provide intuition about the effect of knowledge coverage and model sizes on the complexity of example associations in forgetting. We hope the set of synthetic experiments can inspire more comprehensive study on how the complexity of example associations are affected by various factors in future works.

### v7

@@ Appendix G Towards Interpreting Fine-Grained Associations
Semantic meanings of the $k$-th component in the low-rank approximation of the association matrix $Z$. We perform further analysis into the patterns captured by the $k$-th singular value and singular vectors by identifying the most relevant learned tasks and upstream example domain to the component. For each $k$ and its corresponding component $\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$, we extract top 3 rows with the highest mean (i.e., top 3 relevant learned tasks $T_{i}$). We also extract top 50 columns with highest mean (i.e. top 50 relevant upstream examples) and the domain where these upstream examples are drawn from. For OLMo models, the domains are one of C4, common-crawl, Gutenberg books, Reddit, Science, StackOverFlow, and Wikipedia. We compare the distribution of domains in the top 50 upstream examples, and perform a $z$-test to determine upstream example domain that is significantly more or less forgotten compared to a prior domain distribution of top 50 most forgotten upstream examples (columns with highest mean in $Z$). The results are summarized in Table 10.
@@ Appendix H Effect of Model Training Setups on Forgetting
Low-rank approximations hold at various learning rates, batch sizes, and LoRA fine-tuning. Table 11 summarizes $R^{2}$ of low-rank approximations. Reducing learning rate from the default 2e-6 to 1e-6 causes very minor change of $R^{2}$; Nevertheless, we notice increased $R^{2}$ under a larger learning rate like 5e-6, indicating even simpler associations between the learned tasks and forgotten upstream examples. Under this overly large learning rate, the model suffers substantially more forgetting on all examples. Besides, increasing the batch size (under the same number of parameter update steps) causes an increase in $R^{2}$, indicating simpler associations, possibly due to reduced variance of learned models under a larger batch size. LoRA fine-tuning (Hu et al., 2022) results in simpler associations despite the model forgets less than full parameter fine-tuning.
@@
Replay can change the patterns of forgetting compared to no replay, but the change is mild. We compute sample Pearson correlation coefficient between upstream example forgetting collected under different replay strategies, and more specifically (1) no replay (2) random replay and (3) MF-Offline replay on OLMo-7B and 8 Dolly tasks. We evaluate forgetting of upstream examples held-out from replay. We include correlations between upstream example forgetting under two different random seeds during training as the baseline.
We summarize the results in Table 12. The correlations between no replay and replay are around 0.8, lower than the 0.94 baseline, but are still strongly positive. Replay example selection strategies (random or MF-offline) do not have a strong impact on the correlation. The results imply replay changes patterns of forgetting to a mild degree. Future work can study the limit where the patterns of forgetting experience notable changes.
@@ Appendix I Combining Embedding and Matrix Completion based Prediction
We consider the embedding based approaches have potential to better capture fine-grained knowledge conflicts that leads to forgetting of one example by learning the other (e.g., two documents stating contradictory facts) that is missed by matrix completion approaches that rely on statistics of forgetting.
Appendix J Fine-tuning over Noisy or Adversarial Training Data
(a) OLMo-7B, forgetting over Dolma after fine-tuning over Ultrafeedback with shuffled or flipped preferred responses.
(b) OLMo-7B-Instruct, forgetting over Tulu after fine-tuning over Ultrafeedback with shuffled or flipped preferred responses.
Figure 12: Forgetting after fine-tuning OLMo-7B or OLMo-7B-Instruct on Ultrafeedback dataset with original, shuffled, or flipped preferred responses.
Table 14: Correlations between column mean of $Z$ (in main experiments in Sec. 3, visualized in Figures 2 and 6) and the forgetting occurred on upstream data after fine-tuning on clean, noisy (shuffled) or adversarial (flipped) variants of UltraFeedback.
OLMo-7B OLMo-7B-Instruct
Pearson $\rho$ Spearman $\rho$ Avg. Forgetting Pearson $\rho$ Spearman $\rho$ Avg. Forgetting
Original 0.5884 0.4813 0.0466 0.7382 0.7632 -0.0432
Shuffled 0.5865 0.4591 0.0375 0.4294 0.5624 0.5149
Flipped 0.6535 0.5289 0.0463 0.6516 0.6893 0.1103
In this section, we examine forgetting after fine-tuning on clean, noisy, or adversarial new tasks. We experiment with Ultrafeedback (Cui et al., 2024), an instruction tuning dataset where each instruction is paired with one most preferred response and many rejected responses. For the clean variant, we fine-tune the model over the preferred responses. For the noisy variant, we shuffle the set of the preferred responses so that the instruction and the response mismatch. For the adversarial variant, we fine-tune the model over the rejected responses. We leave preference pair fine-tuning that simultaneously utilize both preferred and rejected responses as future work.
Figure 12 visualizes the forgetting on Dolma or Tulu after fine-tuning OLMo-7B and OLMo-7B instruct, represented in a matrix $Z_{U}\in\mathbb{R}^{3\times N}$. We also include column mean of $Z$ (i.e., upstream example forgetting averaged over all tasks $T_{1..M}$) in the main experiments in Sec. 3 as references. In Table 14, we present the correlations between $Z_{U}$ and the column mean of $Z$.
On OLMo-7B, our results suggests that upstream example forgetting after fine-tuning on rejected responses do not correlates less with the task-averaged upstream example forgetting (column mean of $Z$) compared to fine-tuning on the clean data. On OLMo-7B-Instruct, however, we see a clear drop in the correlation by fine-tuning on the flipped data, very likely because the data contradicts the upstream instruction tuning data. Nevertheless, we notice the correlation with the task-averaged forgetting is still strongly positive. Meanwhile, fine-tuning on shuffled data causes forgetting that is least correlated with the task-averaged upstream example forgetting, indicating more significant change in the patterns of forgetting. Still, we notice a strongly positive correlation.
Appendix K Synthetic Dataset Experiments
@@
To summarize, our analysis with synthetic datasets provide intuition about the effect of knowledge coverage and model sizes on the complexity of example associations in forgetting. We hope the set of synthetic experiments can inspire more comprehensive study on how the complexity of example associations are affected by various factors in future work.